OpenStudy (anonymous):

Find the equation of the line through (4,3) perpendicular to the line 3x+y=7

OpenStudy (anonymous):

so this is going to be similiar to what we did, but for a perpendicular line, the slope is the opposite reciprocal

OpenStudy (anonymous):

it always changes on a perpendicular line?

OpenStudy (anonymous):

i'm not sure what you mean

OpenStudy (anonymous):

if you have two lines that are parallel, they have the same slope if you have two lines that are perpendicular, than their slopes are opposite reciprocals

OpenStudy (anonymous):

like the slope is the opposite reciprocal?

OpenStudy (anonymous):

https://www.purplemath.com/modules/slope2.htm

OpenStudy (anonymous):

yeah, for any two lines to be perpendicular, it means their slope are opposite reciprocals

OpenStudy (anonymous):

ooo ok got it

OpenStudy (anonymous):

you can think of it visually too if you have two lines that are parallel, it means that they never touch, right? so that means that they're at exactly the same angle, like so:  these two lines will always have the same slope  these lines are perpendicular, meaning they're the exact opposite of each other, and they intersect at a 90degree angle so if the slope is 5 for one, the slope for the other will be - 1/5

OpenStudy (anonymous):

ooo ok yea it makes sense

OpenStudy (anonymous):

so using what you learned on the previous problem, can you tell me what the equation is?

OpenStudy (anonymous):

ok so i would put 3x+y=7 then it would be y=3x+7 right?

OpenStudy (anonymous):

keep in mind that you're not just moving it over, you're actually subtracting it from both sides

OpenStudy (anonymous):

o right so it would be y=-3x+7 right?

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

so what's the slope of your perpendicular line?

OpenStudy (anonymous):

ok so now i do the y-y1=m (x-x2) right?

OpenStudy (anonymous):

not quite yet, you don't have your correct slope yet. remember you're looking for the slope of a line that is perpendicular to the line you just gave me

OpenStudy (anonymous):

ooo so i multiply -1 to -3 and +7?

OpenStudy (anonymous):

na you don't even need to do any manipulation to it, just know that the reciprocal of something is the opposite of it so like 2/3 = 3/2 5/4 = 4/5 1/2 = 2/1 5 = 1/5 and if you're taking the opposite, than it's just going to be negative, if your number is positive, and positive if your number is negative so the opposite reiprocal of -9 will be 1/9 and the opposite reciprocal of 7/6 will be -6/7 make sense? so for the line you just gave me, what's its slope and what is its opposite reciprocal?

OpenStudy (anonymous):

so it would be 1/3?

OpenStudy (anonymous):

yup :o

OpenStudy (anonymous):

yay ok so it would be y=1/3+7 right?

OpenStudy (anonymous):

well, you just figured out the slope, now as you said before you have to plug in the two points (4,3) into y-y1 = m (x - x1) you're almost there!!

OpenStudy (anonymous):

ok but when u said 5=1/5 would it be negative too?

OpenStudy (anonymous):

i was just giving examples of reciprocals, without adding in the opposite sign yet the reciprocal of 5 = 1/5, the opposite reciprocal of 5 = -1/5 make sense?

OpenStudy (anonymous):

ook but was i right?

OpenStudy (anonymous):

which part?

OpenStudy (anonymous):

thats 5= 1-/5?

OpenStudy (anonymous):

it depends on what you're looking for.. if you're trying to find the opposite reciprocal, then yes

OpenStudy (anonymous):

the word opposite means that it changes signs (negative to positive, positive to negative) the word reciprocal just means that it flips

OpenStudy (anonymous):

oo ok thnx ok going on to the other problem plug in the two points

OpenStudy (anonymous):

excellent ^.^

OpenStudy (anonymous):

ok i hope im right lol i did y-3=1/3(x-4) right?

OpenStudy (anonymous):

then i did 3(y-3)=x-4

OpenStudy (anonymous):

the first part looks right, but what happened with the y in the second part?

OpenStudy (anonymous):

remember, you want to get the y by itself, so you want to keep that 1/3 over on the right

OpenStudy (anonymous):

are you familiar with distribution?

OpenStudy (anonymous):

yea i am but i thought u leave x were 4 is and 3 on the other side

OpenStudy (anonymous):

what you want to do is distribute the 1/3 to the (x-4)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

just ignore the left side for now, and try to do the distribution

OpenStudy (anonymous):

so u multiply 1/3 with -4?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

and the x as well

OpenStudy (anonymous):

you're literally just distributing the 1/3 into the (x-4)

OpenStudy (anonymous):

u lost me lol

OpenStudy (anonymous):

you said you're familiar with distribution right? in this case, you're distributing the 1/3 into the (x-4) so, you're multiplying both terms by 1/3 like so: (x(1/3) - 4(1/3))

OpenStudy (anonymous):

o yea so it would be1/3x but the other 1 i dont know wat it would be

OpenStudy (anonymous):

if you're multplying 4 by 1/3 it's the same thing as multiplying 4/1 x 1/3, does that make it easier?

OpenStudy (anonymous):

\[\left(\begin{matrix}4 \\ 1\end{matrix}\right) x \left(\begin{matrix}1 \\ 3\end{matrix}\right)\]

OpenStudy (anonymous):

ok so it would be 4/3?

OpenStudy (anonymous):

yup ^.^

OpenStudy (anonymous):

so now, we can look back over at the left side y-3 = 1/3x+4/3

OpenStudy (anonymous):

now, you want to get the y alone, so what should you do?

OpenStudy (anonymous):

ok i would move that negative and put it positve and put it on the other side and add 4/3?

OpenStudy (anonymous):

yup!!! you got it ^.^

OpenStudy (anonymous):

just remember that you can't just add it as it is though, you're dealing with 3/1 + 4/3, so you have to get a common denominator

OpenStudy (anonymous):

ok but how would i add 4/3 to 3?

OpenStudy (anonymous):

\[\left(\begin{matrix}4 \\ 3\end{matrix}\right) + \left(\begin{matrix}3 \\ ?\end{matrix}\right)\]

OpenStudy (anonymous):

oops that ? should be a 1 lol

OpenStudy (anonymous):

lol i was gonna say would it be a 1?

OpenStudy (anonymous):

basically you want to make it so that the two bottom numbers match, so you want it to be something /3 .. whatever you did to the bottom to get that 3, you have to do to the top as well

OpenStudy (anonymous):

\[\left(\begin{matrix}3 \\ 1\end{matrix}\right) = \left(\begin{matrix}9 \\ 3\end{matrix}\right)\], right?

OpenStudy (anonymous):

can you see how i got that?

OpenStudy (anonymous):

no i dont im srry

OpenStudy (anonymous):

so i want to make the bottom match to 4/3 right? i need to do soething to 3/1 so that it has the same bottom ?/3 i multiplied the bottom by 3, to get ?/3 and whatever i did to the bottom, i have to do to the top as well so 3x3=9 /3

OpenStudy (anonymous):

it's kind of hard to explain over text lol

OpenStudy (anonymous):

its ok i get it alittle bit

OpenStudy (anonymous):

but thank you for trying to explain

OpenStudy (anonymous):

when you're adding fractions, the denominators have to be equal.. so if they're not, you have to do something to the bottoms to make them equal,a nd whatever you do to the bottom, you have to do to the top as well

OpenStudy (anonymous):

http://www.mathsisfun.com/fractions_addition.html that explains it pretty well